Write the equation of a quadratic that has zeros of 5 and 1 with F(0)=1. I understand you need to multiply (x-5)(x-1) to get the polynomial, but f(0) wouldnt equal 1. PLEASE HELP ME OUT!!!
Found 2 solutions by nerdybill, stanbon:
Answer by nerdybill(7384)

As you say:
(x-5)(x-1)
x^2 - x - 5x + 5
x^2 - 6x + 5
.
So,
F(x) = x^2 - 6x + 5
.
F(1) = 1^2 - 6(1) + 5
F(1) = 1 - 6 + 5
F(1) = 0
.

Answer by stanbon(75887)

Write the equation of a quadratic that has zeros of 5 and 1 with F(0)=1. I understand you need to multiply (x-5)(x-1) to get the polynomial, but f(0) wouldnt equal 1.
------------------------
Right. That is why you need a variable factor,like "k" below.
---------
f(x) = k(x-5)(x-1)
---
Now you get f(0) = k(-5)(-1) = 1
5k = 1
k = 1/5
------------
EQUATION:
y = (1/5)(x-5)(x-1)
==========================
Cheers,
Stan H.